# ParallelPrompt: Extracting Parallelism from Large Language Model Queries

### v2

@@ Abstract
LLM serving systems typically treat user prompts as monolithic inputs, optimizing inference through decoding tricks or inter-query batching. However, many real-world prompts contain latent semantic parallelism—decomposable structures where subtasks can be executed independently to reduce latency while preserving meaning. We introduce ParallelPrompt, the first benchmark for measuring intra-query parallelism in natural user prompts. Our dataset comprises over 37,000 real-world prompts from public LLM chat logs, each annotated with a structured schema capturing task templates, shared context, and iteration inputs. These schemas are extracted using LLM-assisted prompting with rule-based multilingual validation. To evaluate the benefits of decomposition, we provide an execution suite that benchmarks serial vs. parallel strategies, measuring latency, structural adherence, and semantic fidelity. Our results show that intra-query parallelism can be successfully parsed in over 75% of curated datasets, unlocking up to 5×5$\times$5 × speedups on tasks like translation, comprehension, and comparative analysis, with minimal quality degradation. By releasing this benchmark, curation pipeline, and evaluation suite, we provide the first standardized testbed for studying structure-aware execution in LLM serving pipelines.
@@ 1 Introduction
In this work we focus on natural language prompts that implicitly contain multiple independent subtasks, which we refer to as latent semantic parallelism. For example, asking a model to generate ten short stories will take longer than executing ten generation calls in parallel (see additional examples of parallelizable prompts from public LLM logs in Figure 1). To study this phenomenon at scale, we develop ParallelPrompt, a benchmark of over 37,000 real-world prompts drawn from LMSYS-Chat-1M zheng2023lmsys and WildChat-1M zhao2024wildchat . Our multi-stage pipeline (overview in Figure 2) filters these logs for parallelizable prompts, yielding a surprising 10% of user queries with decomposable structure. High-confidence instances include prompts with clear cues such as explicit lists or numbered items, while medium-confidence ones rely on semantic cues like plural forms or task multiplicity.
Although prior methods have explored related problems in task decomposition, e.g., Skeleton-of-Thought ning2024skeleton , Tree-of-Problems zebaze2024tree , and LLMCompiler kim2024llm , these approaches target synthetic or tool-based settings. In contrast, ParallelPrompt provides naturally occurring, free-form prompts submitted by real users. Interestingly, we find that prior task decomposition methods fail on many of the prompts in our benchmark—highlighting the importance of evaluating on real user inputs.
[…]
@@ 3.1 Benchmark Design and Curation
[…]
@@ 3.2 Validation Methodology and Quality Control
Our validation process applies several key constraints to ensure schema integrity, including mutual exclusivity between data and count fields, template-placeholder compatibility, and minimum parallelism thresholds. Common failure patterns and their frequencies are analyzed in Appendix B.3, providing insights into the limitations of current schema extraction approaches.
@@ 3.3 Dataset Statistics and Research Implications
Scale and Prevalence. From 358,000 inspected prompts, we validated 37,070 as parallelizable (16,721 from LMSYS and 20,349 from WildChat), representing a 10.3% yield. This finding challenges the assumption that intra-query parallelism is rare, revealing instead that it constitutes a significant mode of user interaction with large language models.
@@ 4.1 Evaluation Methodology
Our evaluation pipeline operates in two phases: (1) schema extraction using Claude 3.5, and (2) execution using GPT-4-1106-preview via the OpenAI API. This separation mirrors real-world deployments where decomposition and execution might be handled by different components. Implementation details, including thread management and timeout handling, are provided in Appendix F.1.
@@ 4.2 Performance Results
Keyword Extraction 2.38 3.23 2.54$\times$ 1.36$\times$
Reading Comprehension 3.49 10.27 5.72$\times$ 2.94$\times$
Repeated Generation 3.79 9.51 4.39$\times$ 2.50$\times$
Repeated Generation (E2E) 4.88 9.51 3.41$\times$ 1.70$\times$
Latency Reduction. As shown in Table 1, parallel execution yields substantial performance improvements across all evaluated categories. Reading Comprehension tasks show the highest raw speedup (5.72$\times$), with Repeated Generation (4.39$\times$) and Keyword Extraction (2.54$\times$) also demonstrating significant gains. Importantly, even when including schema extraction overhead in end-to-end measurements, Repeated Generation tasks still achieve a 3.41$\times$ speedup—highlighting the practical value of intra-query parallelism in realistic deployment scenarios.
@@
Comparative Method Performance. Our quantitative comparison with SoT and ToP, detailed in Appendix I, reveals the limitations of specialized decomposition approaches when faced with diverse real-world prompts. While these methods perform well within their designed domains (outline expansion for SoT, recursive problem decomposition for ToP), they struggle with the varied parallelization patterns present in our benchmark.
@@ 4.4 Practical Implications
  • Significant Latency Reductions: Even accounting for schema extraction overhead and quality normalization, parallelization yields 1.4-3$\times$ speedups across diverse task types.
@@ 5.1 Beyond Structural Decomposition
The evaluation results in Section 4 highlight an important distinction: structurally decomposable prompts don’t always benefit from parallel execution. This gap between structural identification and effective parallelization appears most clearly in prompts with subtle dependencies: for example, consider — "Compare each paragraph to the previous one and analyze thematic shifts.". While our extraction pipeline identifies structural parallelism (multiple paragraphs to analyze), the semantic dependencies between subtasks make parallel execution ineffective. This pattern of dependency blindness, analyzed in detail in Appendix D.2, affects even high-confidence validations.
@@
Despite these challenges, our manual inspection confirms that over 75% of structurally validated prompts remain suitable for effective parallelization, suggesting broad applicability to real-world LLM workloads. This finding points to a productive middle ground between entirely serial execution and naive decomposition: task-adaptive strategies that recognize when parallelism will preserve semantic integrity.
@@ 5.2 Linguistic and Demographic Biases
These biases reflect both the demographic composition of early LLM adopters and structural assumptions in our extraction approach. While our benchmark achieves substantial multilingual coverage, future work should prioritize language-aware extraction methods that account for diverse syntactic conventions and prompt formulation strategies.
Beyond linguistic representation, the source datasets (WildChat-1M and LMSYS-chat-1M) capture interaction patterns from early adopters who may differ from broader populations in their technical sophistication and prompt complexity. While our sampling approach ($\approx$5,500 manual inspections) provides statistical power for major patterns, conclusions about underrepresented languages and long-tailed categories should be interpreted cautiously.
@@ 5.4 Structure-Aware Execution in Practice
The $\approx$23failure rate in our manual inspection is comparable to widely-used datasets in other domains (beyer2020we, ; ribeiro2020beyond, ; paullada2021data, ; northcutt2021pervasive, ), suggesting that despite its limitations, structure-aware execution could be a valuable optimization for most real-world LLM workloads.
@@ E.2 Cross-Linguistic Examples
Çompara estos tres filósofos en términos de sus ideas principales…"missing
@@ F.1.1 Schema Extraction Phase
We use Claude 3.5 (AWS Bedrock) as the schema extraction engine. Given a user prompt, Claude generates a structured JSON object capturing the decomposition schema, including task template, context, data items or counts, and category. This structured output is validated using a combination of rule-based checks and confidence scoring, as described in Appendix B.
@@ F.3 Diversity Enforcement for Repeated Generation
This simple heuristic reduces redundancy in large-scale generation tasks and ensures that parallel outputs are meaningfully distinct, particularly important for creative generation tasks where the model might otherwise produce similar variations when executed in parallel without context awareness. Our analysis shows this constraint increases output diversity by 37% based on semantic similarity measures.
@@ G.1 Performance Metrics
    Normalized Speedup=Raw Speedup×Parallel Token CountSerial Token $\text{Normalized Speedup}=\text{Raw Speedup}\times\frac{\text{Parallel Token Count}}{\text{Serial Token Count}}$ (1)

### v3

@@ Abstract
LLM serving systems typically treat user prompts as monolithic inputs, optimizing inference through decoding tricks or inter-query batching. However, many real-world prompts contain latent semantic parallelism—decomposable structures where subtasks can be executed independently to reduce latency while preserving meaning. We introduce ParallelPrompt, the first benchmark for measuring intra-query parallelism in natural user prompts. Our dataset comprises over 37,000 real-world prompts from public LLM chat logs, each annotated with a structured schema capturing task templates, shared context, and iteration inputs. These schemas are extracted using LLM-assisted prompting with rule-based multilingual validation. To evaluate the benefits of decomposition, we provide an execution suite that benchmarks serial vs. parallel strategies, measuring latency, structural adherence, and semantic fidelity. Our results show that intra-query parallelism can be successfully parsed in over 75% of curated datasets, unlocking up to $5\times$ speedups on tasks like translation, comprehension, and comparative analysis, with minimal quality degradation. By releasing this benchmark, curation pipeline, and evaluation suite, we provide the first standardized testbed for studying structure-aware execution in LLM serving pipelines.
@@ 1 Introduction
In this work we focus on natural language prompts that implicitly contain multiple independent subtasks, which we refer to as latent semantic parallelism. For example, asking a model to generate ten short stories will take longer than executing ten generation calls in parallel (see examples of such prompts from public LLM logs in Figure 1). To study this phenomenon at scale, we develop ParallelPrompt, a benchmark of over 37,000 real-world prompts drawn from LMSYS-Chat-1M zheng2023lmsys and WildChat-1M zhao2024wildchat . Our multi-stage pipeline (overview in Figure 2) filters these logs for parallelizable prompts, yielding a surprising 10% of user queries with decomposable structure. This validation rate represents a conservative, high-precision estimate that translates to practical impact: in production deployments handling billions of monthly queries altman2025chatgpt ; pichai2025aioverviews ; google2024speculative , this yields millions of optimization opportunities with latency improvements. For context, even widely adopted optimizations like speculative decoding leviathan2023spec ; chen2023accelerating and KV-caching kwon2023efficient only benefit certain query types, yet are considered valuable production techniques. Perhaps more importantly, intra-query parallelism complements rather than competes with existing optimizations, offering further gains within individual queries.
Although prior methods have explored related problems in task decomposition, e.g., Skeleton-of-Thought ning2024skeleton , Tree-of-Problems zebaze2024tree , and LLMCompiler kim2024llm , these approaches target synthetic scenarios or predefined schemas. Interestingly, we find that prior task decomposition methods fail on many of the prompts in our benchmark. This limitation stems from fundamental architectural constraints: prior methods make strong assumptions about task structure that don’t generalize to the diverse parallelization patterns present in real user prompts. In contrast, our addresses naturally occurring user prompts where parallelizable structure must be discovered rather than assumed.
ParallelPrompt enables both method and system evaluation. Our results show that naive serial execution is often suboptimal, with simple parallelization yielding $3{\times}$$5{\times}$ latency improvements on tasks like translation and comprehension. We also surface failure cases like dependency blindness, highlighting open challenges for future work. Overall, we make the following contributions:
@@ 3.1 Benchmark Design and Curation
Our validation pipeline explicitly prioritizes precision over recall, where precision represents the fraction of validated prompts that can be executed in parallel without semantic degradation, and recall represents the fraction of all inherently parallelizable prompts that our pipeline successfully identifies. This conservative approach ensures benchmark quality while acknowledging that true parallelism rates could be higher with improved extraction methods and more so for specialized domains.
@@
Multilingual Coverage and Extraction Challenges. ParallelPrompt captures parallelizable prompts across at least 11 languages. While English dominates ($\approx\qty$84), Figure 9 (in the appendix) reveals substantial representation in Chinese (5.6%), Russian (6.3%), and other languages. Importantly, validation success rates vary significantly across languages: European languages show higher structural reliability (55–63% high-confidence), while East Asian languages like Chinese and Japanese exhibit lower rates (28–34%). This disparity reflects both linguistic structural differences and extraction method biases, as explored in Appendix E.1.
Rather than presenting artificially balanced data that would obscure real-world deployment challenges, we, however, preserved the authentic distributional patterns of user traffic (84% English) to expose genuine multilingual fragilities that require addressing. This positions ParallelPrompt as the first benchmark to systematically analyze multilingual decomposition challenges, providing detailed failure pattern analysis (Appendix E.1) to guide development of language-aware extraction methods.
@@ 3.2 Validation Methodology and Quality Control
[…]
The validation process applies several key constraints to ensure schema integrity, including mutual exclusivity between data and count fields, template-placeholder compatibility, and minimum parallelism thresholds. Common failure patterns (analyzed in Appendix B.3) include template-placeholder mismatches, mutual exclusivity violations between data and count fields, and insufficient parallelism thresholds, providing insights into the limitations of current schema extraction approaches.
@@ 3.3 Dataset Statistics and Research Implications
Scale and Prevalence. From 358,000 inspected prompts, we validated 37,070 as parallelizable (16,721 from LMSYS and 20,349 from WildChat), representing a 10.3% yield. This finding challenges the assumption that intra-query parallelism is rare, revealing instead that it constitutes a significant mode of user interaction with LLMs.
@@
3.4 Benchmark Applications
ParallelPrompt enables three primary evaluation scenarios: (1) Schema extraction research, evaluating how well different models or prompting strategies identify parallelizable structure; (2) Decomposition execution strategies, through systematic latency-quality tradeoff evaluation across diverse task categories; and (3) Model benchmarking, assessing how different LLMs perform under parallel versus serial execution. The task-agnostic schema format and modular C++ infrastructure support any structure-aware execution pipeline, providing a standardized testbed for advancing structure-aware LLM serving research beyond the specific method evaluated here.
@@ 4.1 Evaluation Methodology
Our evaluation pipeline operates in two phases: (1) schema extraction using Claude 3.5, and (2) execution using GPT-4-1106-preview via the OpenAI API. For schema extraction, we evaluated multiple models including Claude-3.7 Sonnet (too powerful, over-extracted with excessive decomposition), GPT-4o-mini (too weak, under-performed with high error rates), and several open-source models via Together.ai API before selecting Claude 3.5 for its optimal precision-coverage tradeoff. Implementation details, including thread management and timeout handling, are provided in Appendix F.1.
@@
The normalized speedup metric addresses a systematic bias where LLMs processing multiple items sequentially often truncate individual responses due to internal "length budgeting" behavior son2024multi ; cheng2023batch ; zheng2023response . For example, requesting one story might yield 500 words, while requesting ten stories results in  50 words each as the model attempts to fit everything into similar total response lengths. Parallel execution avoids this by allocating full context to each subtask, and our metric adjusts for these length differences to enable fair efficiency comparisons.
@@ 4.2 Performance Results
[…]
@@
Comparative Method Performance. Our quantitative comparison with SoT and ToP, detailed in Appendix I, reveals the limitations of specialized decomposition approaches when faced with diverse real-world prompts. SoT’s two-phase outline-then-expand approach makes assumptions about task structure that don’t generalize to extraction tasks, while ToP’s recursive-problem-focused decomposition adds unnecessary overhead for flat, independent tasks; hence, they struggle with the varied parallelization patterns present in our benchmark.
@@ 4.4 Practical Implications
[…]
@@ 5.1 Beyond Structural Decomposition
The evaluation results in Section 4 highlight an important distinction: structurally decomposable prompts don’t always benefit from parallel execution. This gap between structural identification and effective parallelization appears most clearly in prompts with subtle dependencies: for example, consider — "Compare each paragraph to the previous one and analyze thematic shifts." While our extraction pipeline identifies structural parallelism (multiple paragraphs to analyze), the semantic dependencies between subtasks make parallel execution ineffective. This pattern of dependency blindness, analyzed in detail in Appendix D.2, affects even high-confidence validations.
@@
Dependency blindness affects $\approx$25% of structurally valid prompts, manifesting in three primary types: sequential dependencies (comparative analysis requiring previous context), semantic dependencies (narrative coherence), and constraint dependencies (shared formatting requirements). These challenges suggest a productive middle ground between entirely serial execution and naive decomposition: task-adaptive strategies that recognize when parallelism will preserve semantic integrity.
@@ 5.2 Linguistic and Demographic Biases
These biases reflect both the demographic composition of early LLM adopters and structural assumptions in our extraction approach that may differ from broader populations in technical sophistication and prompt complexity. While our benchmark achieves substantial multilingual coverage, future work should prioritize language-aware extraction methods that account for diverse syntactic conventions and prompt formulation strategies.
Beyond linguistic representation, the source datasets (WildChat-1M and LMSYS-chat-1M) capture interaction patterns from early adopters who may differ from broader populations in their technical sophistication and prompt complexity. While our sampling approach ($\approx$5,500 manual inspections) provides statistical power for major patterns, we view these biases as exposing genuine multilingual challenges in LLM deployments that need addressing, rather than fundamental limitations of the approach.
@@ 5.4 Structure-Aware Execution in Practice
The $\approx$23% failure rate in our manual inspection is comparable to widely-used datasets in other domains (beyer2020we, ; ribeiro2020beyond, ; paullada2021data, ; northcutt2021pervasive, ), suggesting that despite its limitations, structure-aware execution could be a valuable optimization for most real-world LLM workloads.
@@ 6 Conclusion
Acknowledgements
We thank Daniel Fried for the helpful reviews and comments during the development of this work.
@@ E.2 Cross-Linguistic Examples
Çompara estos tres filósofos en términos de sus ideas principales…"
@@ F.1.1 Schema Extraction Phase
We use Claude 3.5 Haiku (AWS Bedrock) as the schema extraction engine. Given a user prompt, Claude generates a structured JSON object capturing the decomposition schema, including task template, context, data items or counts, and category. This structured output is validated using a combination of rule-based checks and confidence scoring, as described in Appendix B.
@@ F.3 Diversity Enforcement for Repeated Generation
This simple heuristic reduces redundancy in large-scale generation tasks and ensures that parallel outputs are meaningfully distinct, particularly important for creative generation tasks where the model might otherwise produce similar variations when executed in parallel without context awareness. Our skimmed analysis shows this constraint increases output diversity by 57% based on semantic similarity measures.
@@ G.1 Performance Metrics
    $\text{Normalized Speedup}=\text{Raw Speedup}\times\frac{\text{Parallel Token Count}}{\text{Serial Token Count}}$ (1)